# An Autonomous Vision-Based Algorithm for Interplanetary Navigation

### v2

@@ 5.2.3 Simulation Settings
The VBN filter proposed in this work is tested on an interplanetary high-fidelity ballistic trajectory between Earth–Mars [49]. The dynamics of the reference true trajectory include the SRP perturbations, the main attractor acceleration, third-body accelerations due to all the planets in the Solar System, and relativistic perturbations. Note that the dynamic model selected for the filter (Eq. 24) is a lower-fidelity one, implying that the unmodeled accelerations are captured by the GM processes. Fig. 5 shows the analyzed leg of the nominal probe trajectory. Note that this is an improvement with respect to Andreis et al. [9] where simpler dynamics models are used in the filter and in the truth simulation.
@@
Table 5: Onboard Camera Characteristics
FoV [deg] F [-] T [ms] Image size [px] f [mm] Qe×$\times$ Tlens $\sigma_{d}$ [px] $n_{\rm{CR}}$ SEA
20 2.2 400 1024$\times$ 1024 40 0.49 0.5 3 20
Figs. 5(a) and 5(b) are two of the rendered deep-space sky images adopted in the filtering procedure.
(a) Image taken during Earth observation in LIS mode
(b) Image taken during Mars observation in LIS mode
Figure 6: Generated sky-field images. The planet is outlined with a green marker and identified stars with a white marker
For what concerns the IP, the parameters adopted in the thresholding step and the RANSAC are detailed in Table 6. In detail, $k_{\textrm{thr}}$ is the tuning parameter of the Niblack’s method, $n_{R}$ is the total number of samples considered, $m_{R}$ is the number of centroids for each sample, and $t_{R}$ is the threshold value considered for data fitting. More details about the RANSAC procedure can be found in Andreis et al. [24].
Table 6: Parameters of the RANSAC and thresholding procedures
$k_{\textrm{thr}}$ [-] $n_{R}$ [-] $m_{R}$ [-] $t_{R}$ [arcsec]
10 30 3 100
@@
$\sigma_{{r}}$ [km] $\sigma_{{v}}$ [km/s] $\sigma_{\eta_{\textrm{SRP}}}$[km/$s^{2}$] $\sigma_{\eta_{\textrm{R}}}$ [km/$s^{2}$]
$10^{4}$ $10^{-1}$ $10^{-11}$ $10^{-11}$
Note that the values are selected following a conservative approach, taking into account that in deep space the initial position and velocity are usually known with an accuracy better than $10^{4}$ km and $0.1$ km/s, respectively. However, even if the performance of the IP procedure degrades by increasing the probe position uncertainty, this has been tested to work up to $\sigma_{r}=10^{6}$ km with a success rate higher than 90% in planet identification. Then, the performance of the IP worsens to 70% when $\sigma_{r}=10^{7}$ km (see Sec. 5.1).For what concerns the OD, in Andreis et al. [9] the algorithm is tested to work up to $\sigma_{r}=10^{7}$ km. Over this value, the OD algorithm is not able to select the optimal targets.
[…]
@@ 5.2.4 Filter Performance
To evaluate the filter performance, a Monte Carlo simulation of 100 samples is performed. The lower number of simulations is due to the need to simulate long propagation arcs and the large number of images for every Monte Carlo sample. Figures 7 and 8 show the position and velocity error profiles and $3\sigma$ covariance bounds in the J2000 ecliptic reference frame on the studied trajectory leg. The sample error profile is displayed with blue solid lines, whereas the orange solid lines and the dashed ones define the $3\sigma$ covariance bounds of the samples and the filter, respectively.
Figure 7: Estimated errors for each position component with related $3\sigma$ bounds
@@
  1. MC2:
    Only a correction of the light-time effect acting on the position of the planets. The measurement model is defined by Eqs.  49 and 50;
  2. MC3:
    Only a correction of the light-aberration effect acting on the position of the planets. Therefore, $\Delta t=0$ and $\delta\Delta t=0$ in Eqs. 37 and 47;
  3. MC4:
    No correction of the light effects acting on the position of the planets. The measurement model is defined by Eqs.  49 and 50 with $\Delta t=0$ and $\delta\Delta t=0$ in Eqs.37 and 47;
  4. MC5:
    A correction of the light-time effect acting on the position of the planets. Hence, the measurement model is defined by Eqs.  49 and 50. Furthermore, the stellar light aberration is not corrected in the IP.
Note that in the first four Monte Carlo simulations (i.e., MC1, MC2, MC3, and MC4), the stellar aberration correction is always implemented in the IP (see Sec. 3.2) despite what is implemented in the measurement model. This assumption was performed to generalize the analysis to the case in which an attitude estimation already compensated for light aberration is provided to the filter either by the attitude determination and control system or directly from a star tracker. Moreover, an uncorrected attitude would induce a constant bias in the filter, leading to an unfair comparison among the different implementations. Therefore, the analysis is focused on the potential implications if the planets locations remain uncorrected for light effects. Conversely, the last MC study is relevant when the same optical camera exploited for navigation serves for attitude determination. In this case, the option to not compensate for stellar aberration exists, potentially yielding an aberrated attitude value. Hence, this scenario explores situations where neither stellar nor planetary aberrations are corrected.
@@ 4. MC5:
The root-mean-square error (RMSE) is chosen as the performance index to measure the estimation accuracy of the different models for the selected dataset as it captures effects due to mean (e.g., biases) and covariance of the estimated solution. The latter comprises 100 initial state vectors and 740 images for simulation representing the scenarios observed during the interplanetary transfer. The filter and camera settings are unchanged with respect to MC1 for the sake of comparison.
(a) Position RMSE
(b) Velocity RMSE
Figure 9: Comparison of the performance of the filter under various light effects
@@
Fig. 9 shows the results of the analysis. The most significant deviation from the nominal value, represented by MC1 and whose components are shown in Figs. 7 and 8, occurs when both light effects corrections are not taken into account in the filter (MC4). Between the two effects, the greatest deviation from the nominal value is obtained when the light-aberration effect remains uncompensated (MC2), compared to the scenario where the light-time effect is corrected (MC3). When the light aberration correction is not applied on both the stars and planets centroids, the probe state estimation is affected by a negligible deviation (MC5). This occurs since, being the attitude adopted in the measurement model equation determined directly from the image, the light aberration affects similarly the measurement model and the external observation. Therefore, is evident that when the same optical sensor is adopted for attitude determination and navigation, the light-aberration correction may not be necessary, saving computational time.
Eventually, validation of the VBN filter performance can be achieved by comparing results with those obtained from a concurrent in-flight demonstration of celestial triangulation with distant planets [51]. This validation repeats the simulation on the ballistic Martian corridor by setting position and velocity uncertainties to values matching those utilized in the in-flight testing (i.e., $\sigma_{r}=10^{5}$ km and $\sigma_{v}=1$ km/s). During the demonstration, observations of Jupiter, Mars, and Saturn were made 8 times, once, and 7 times respectively, with measurement errors set at 0.5 px, 0.75 px, and 0.25 px. To accurately model the measurement error covariance matrix in the implemented filter, unit measurement uncertainty is set to $\sigma_{\rm{str}}=0.2=0.4*36.6/67$, considering an weighted average value for the measurement error in pixels (i.e., 0.4 pixel) multiplied by the camera iFoV specified in Krause et al. [51] (i.e., 36.6 arcsec) and divided by the camera iFoV employed in the current simulation (i.e., 67 arcsec). This is done to compare measurements with the same angular error despite the camera carried on board and eventual calibration residual errors. Additionally, the implemented filter assumes to acquire only 16 observations throughout the trajectory to match the number of images adopted in the in-flight testing. At the end of the trajectory, the filter estimates position and velocity with a $3\sigma$ accuracy of 12000 km and 2.15 m/s respectively. The order of magnitude of the obtained residuals is consistent with the one found in Krause et al. [51] (i.e., 60000 km and 5 m/s).
@@ 6 Conclusion
This paper develops an autonomous vision-based navigation algorithm for interplanetary transfer with an application for CubeSats missions. An extended Kalman filter adopting the planet location as external observation is chosen by considering the limited processing capabilities of a standard miniaturized processor. Moreover, the measurements exploited for the estimation correction are directly extracted from images generated with a deep-space rendering engine. This procedure allows for obtaining a more faithful value of the measurement error and its influence on the filter solution. At the end of the Earth–Mars trajectory, the filter estimates the spacecraft position and velocity with an accuracy of 2000 km and 0.5 m/s, respectively. Future analysis should examine the performance of the filter over low-thrust trajectories, which are desirable in CubeSats applications. Moreover, this work determines the probe attitude from deep-space images with state-of-the-art star identification and attitude determination algorithms. Integrating the attitude filter in the proposed vision-based navigation algorithm is object of future investigation. Eventually, to validate the vision-based navigation applicability, it will be tested through hardware-in-the-loop simulations. Preliminary results have already been obtained through, first, the validation of the orbit determination algorithm with processor-in-the-loop simulations [9] and, second, with the validation of the image processing pipeline with the optical facility in the loop [52, 53, 54, 55]. More details about the validation and verification procedure carried on to increase the TRL of the navigation algorithm within the EXTREMA project are reported in Di Domenico et al. [56] and Di Domenico et al. [57].
@@ Acknowledgments
Appendix
Appendix A Useful matrices to compute derivatives of the measurement model
$\bm{R}_{\rm sc}=\left(\bm{I}_{3x3}+2\,{}^{\mathcal{}}\bm{\beta}_{\rm{sc}}{}^{\mathcal{}}\bm{l}_{\rm{pl/sc}}^{\top}-{}^{\mathcal{}}\bm{l}_{\rm{pl/sc}}^{\top}{}^{\mathcal{}}\bm{\beta}_{\rm{sc}}{\color[rgb]{0,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,0}\pgfsys@color@gray@stroke{0}\pgfsys@color@gray@fill{0}I_{3x3}}-{}^{\mathcal{}}\bm{l}_{\rm{pl/sc}}{}^{\mathcal{}}\bm{\beta}_{\rm{sc}}^{\top}\right)\left(\frac{\bm{I}_{3x3}}{\left|\left|\left(\bm{r}_{\rm{pl}}(t-\Delta t)-\bm{r}(t)\right)^{\top}\left(\bm{r}_{\rm{pl}}(t-\Delta t)-\bm{r}(t)\right)\right|\right|}+\right.\\ \left.-\frac{\left(\bm{r}_{\rm{pl}}(t-\Delta t)-\bm{r}(t)\right)\left(\bm{r}_{\rm{pl}}(t-\Delta t)-\bm{r}(t)\right)^{\top}}{\left|\left|\left(\bm{r}_{\rm{pl}}(t-\Delta t)-\bm{r}(t)\right)^{\top}\left(\bm{r}_{\rm{pl}}(t-\Delta t)-\bm{r}(t)\right)\right|\right|^{3}}\right)\left(\frac{\bm{v}_{\rm{pl}}}{c^{2}-\bm{v}_{\rm{pl}}(t)^{\top}\bm{v}_{\rm{pl}}(t)}\right.\\ \left.\Biggl{(}\frac{\left(c^{2}-\bm{v}_{\rm{pl}}(t)^{\top}\bm{v}_{\rm{pl}}(t)\right)\bm{r}_{\rm{pl/sc}}(t)^{\top}+\bm{r}_{\rm{pl/sc}}(t)^{\top}\bm{v}_{\rm{pl}}(t)\bm{v}_{\rm{pl}}(t)^{\top}}{\sqrt{\bm{r}_{\rm{pl/sc}}(t)^{\top}\bm{r}_{\rm{pl/sc}}(t)\left(c^{2}-\bm{v}_{\rm{pl}}(t)^{\top}\bm{v}_{\rm{pl}}(t)\right)+\left(\bm{r}_{\rm{pl/sc}}(t)^{\top}\bm{v}_{\rm{pl}}(t)\right)^{2}}}-\bm{v}_{\rm{pl}}^{\top}(t)\Biggr{)}-\bm{I}_{3x3}\right)$ (57)
$\bm{V}_{\rm sc}=\frac{1}{c}\left({}^{\mathcal{}}\bm{l}_{\rm{pl/sc}}^{\top}{}^{\mathcal{}}\bm{l}_{\rm{pl/sc}}{\color[rgb]{0,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,0}\pgfsys@color@gray@stroke{0}\pgfsys@color@gray@fill{0}I_{3x3}}-{}^{\mathcal{}}\bm{l}_{\rm{pl/sc}}{}^{\mathcal{}}\bm{l}_{\rm{pl/sc}}^{\top}\right)$ (58)
$\bm{Q}_{\rm att}=2\left[{}^{\mathcal{}}\bm{l}_{\rm{pl/sc}}^{\rm{aberr}}\right]^{\wedge}$ (59)
$\bm{R}_{\rm pl}=\left(\bm{I}_{3x3}+2\,{}^{\mathcal{}}\bm{\beta}_{\rm{sc}}{}^{\mathcal{}}\bm{l}_{\rm{pl/sc}}^{\top}-{}^{\mathcal{}}\bm{l}_{\rm{pl/sc}}^{\top}{}^{\mathcal{}}\bm{\beta}_{\rm{sc}}{\color[rgb]{0,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,0}\pgfsys@color@gray@stroke{0}\pgfsys@color@gray@fill{0}I_{3x3}}-{}^{\mathcal{}}\bm{l}_{\rm{pl/sc}}{}^{\mathcal{}}\bm{\beta}_{\rm{sc}}^{\top}\right)\left(\frac{\bm{I}_{3x3}}{\left|\left|\left(\bm{r}_{\rm{pl}}(t-\Delta t)-\bm{r}(t)\right)^{\top}\left(\bm{r}_{\rm{pl}}(t-\Delta t)-\bm{r}(t)\right)\right|\right|}+\right.\\ \left.-\frac{\left(\bm{r}_{\rm{pl}}(t-\Delta t)-\bm{r}(t)\right)\left(\bm{r}_{\rm{pl}}(t-\Delta t)-\bm{r}(t)\right)^{\top}}{\left|\left|\left(\bm{r}_{\rm{pl}}(t-\Delta t)-\bm{r}(t)\right)^{\top}\left(\bm{r}_{\rm{pl}}(t-\Delta t)-\bm{r}(t)\right)\right|\right|^{3}}\right)\Biggl{(}\bm{I}_{3x3}+\\ -\frac{\bm{v}_{\rm{pl}}}{c^{2}-\bm{v}_{\rm{pl}}(t)^{\top}\bm{v}_{\rm{pl}}(t)}\Biggl{(}\frac{\left(c^{2}-\bm{v}_{\rm{pl}}(t)^{\top}\bm{v}_{\rm{pl}}(t)\right)\bm{r}_{\rm{pl/sc}}(t)^{\top}+\bm{r}_{\rm{pl/sc}}(t)^{\top}\bm{v}_{\rm{pl}}(t)\bm{v}_{\rm{pl}}(t)^{\top}}{\sqrt{\bm{r}_{\rm{pl/sc}}(t)^{\top}\bm{r}_{\rm{pl/sc}}(t)\left(c^{2}-\bm{v}_{\rm{pl}}(t)^{\top}\bm{v}_{\rm{pl}}(t)\right)+\left(\bm{r}_{\rm{pl/sc}}(t)^{\top}\bm{v}_{\rm{pl}}(t)\right)^{2}}}-\bm{v}_{\rm{pl}}^{\top}(t)\Biggr{)}\Biggr{)}$ (60)
$\bm{V}_{\rm pl}=-\left(\bm{I}_{3x3}+2\,{}^{\mathcal{}}\bm{\beta}_{\rm{sc}}{}^{\mathcal{}}\bm{l}_{\rm{pl/sc}}^{\top}-{}^{\mathcal{}}\bm{l}_{\rm{pl/sc}}^{\top}{}^{\mathcal{}}\bm{\beta}_{\rm{sc}}{\color[rgb]{0,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,0}\pgfsys@color@gray@stroke{0}\pgfsys@color@gray@fill{0}I_{3x3}}-{}^{\mathcal{}}\bm{l}_{\rm{pl/sc}}{}^{\mathcal{}}\bm{\beta}_{\rm{sc}}^{\top}\right)\left(\frac{\bm{I}_{3x3}}{\left|\left|\left(\bm{r}_{\rm{pl}}(t-\Delta t)-\bm{r}(t)\right)^{\top}\left(\bm{r}_{\rm{pl}}(t-\Delta t)-\bm{r}(t)\right)\right|\right|}+\right.\\ \left.-\frac{\left(\bm{r}_{\rm{pl}}(t-\Delta t)-\bm{r}(t)\right)\left(\bm{r}_{\rm{pl}}(t-\Delta t)-\bm{r}(t)\right)^{\top}}{\left|\left|\left(\bm{r}_{\rm{pl}}(t-\Delta t)-\bm{r}(t)\right)^{\top}\left(\bm{r}_{\rm{pl}}(t-\Delta t)-\bm{r}(t)\right)\right|\right|^{3}}\right)\Biggl{(}\Delta tI_{3x3}+\\ +\frac{\bm{v}_{\rm{pl}}}{c^{2}-\bm{v}_{\rm{pl}}(t)^{\top}\bm{v}_{\rm{pl}}(t)}\Biggl{(}\frac{\bm{r}_{\rm{pl/sc}}(t)^{\top}\bm{v}_{\rm{pl}}(t)\bm{r}_{\rm{pl/sc}}(t)^{\top}-\bm{r}_{\rm{pl/sc}}(t)^{\top}\bm{r}_{\rm{pl/sc}}(t)\bm{v}_{\rm{pl}}(t)^{\top}}{\sqrt{\bm{r}_{\rm{pl/sc}}(t)^{\top}\bm{r}_{\rm{pl/sc}}(t)\left(c^{2}-\bm{v}_{\rm{pl}}(t)^{\top}\bm{v}_{\rm{pl}}(t)\right)+\left(\bm{r}_{\rm{pl/sc}}(t)^{\top}\bm{v}_{\rm{pl}}(t)\right)^{2}}}+\\ +\frac{2\bm{v}_{\rm{pl}}(t)^{\top}}{c^{2}-\bm{v}_{\rm{pl}}(t)^{\top}\bm{v}_{\rm{pl}}(t)}-\bm{r}_{\rm{pl/sc}}(t)^{\top}\Biggr{)}\Biggr{)}$ (61)

### v3

@@ 5.2.2 Simulation Settings
The VBN filter proposed in this work is tested on an interplanetary high-fidelity ballistic trajectory between Earth–Mars [44]. The dynamics of the reference true trajectory include the SRP perturbations, the main attractor acceleration, third-body accelerations due to all the planets in the Solar System, and relativistic perturbations. Note that the dynamic model selected for the filter (Eq. LABEL:eq:f) is a lower-fidelity one, implying that the unmodeled accelerations are captured by the GM processes. Fig. 5 shows the analyzed leg of the nominal probe trajectory.
@@
Table 4: Onboard Camera Characteristics.
FoV [deg] F [-] T [ms] Image size [px] f [mm] Qe×$\times$ Tlens $\sigma_{d}$ [px] $n_{\rm{CR}}$ SAE
20 2.2 400 1024$\times$ 1024 40 0.49 0.5 1 20
Figs. 6(a) and 6(b) are two of the rendered deep-space sky images adopted in the filtering procedure.
(a) Image taken during first planet observation
(b) Image taken during second planet observation
Figure 6: Generated Sky-field Images
@@
$\sigma_{\textrm{r}}$ [km] $\sigma_{\textrm{v}}$ [km/s] $\sigma_{\textrm{SRP}}$ [km/$s^{2}$] $\sigma_{\textrm{R}}$ [km/$s^{2}$]
$10^{4}$ $10^{-1}$ $10^{-10}$ $10^{-10}$
Note that the values are selected following a conservative approach, taking into account that in deep space the initial position and velocity are usually known with an accuracy better than $10^{4}$ km and $0.1$ km/s, respectively. However, even if the performance of the IP procedure degrades by increasing the probe position uncertainty, this has been tested to work up to $\sigma_{r}=10^{6}$ km with a success rate higher than 90% in planet detection. Then, the performance of the IP worsens to 70% when $\sigma_{r}=10^{7}$ km (see Sec. 5.1) under the assumption of Gaussian errors in the measurement. For what concerns the OD, in [9] the algorithm has been tested to work up to $\sigma_{r}=10^{7}$ km. Over this value, the OD algorithm is not able to select the optimal targets. Moreover, the standard deviation of the measurement error is set to $\sigma_{\rm{str}}=0.1$ px, considering the results of the Monte Carlo runs in the extraction of the planet centroid reported in Sec. 5.1. Eventually, only planets whose apparent magnitude is lower than 7 and whose SAE is greater than 20° are assumed to be visible by the camera, therefore, they are the only ones considered available for the optimal beacon selection process.
@@ 5.2.3 Filter Performance
A Monte Carlo simulation of 50 samples is performed. The initial probe state vector is perturbed by applying the $3\sigma$ standard deviation rule:
$\bm{x}_{0}=\tilde{\bm{x}}_{0}+3\sqrt{\bm{P}_{0}}\bm{k}$ (45)
where $\tilde{\bm{x}}_{0}$ is the probe nominal state, $\bm{k}$ is a random vector with values within $[-1;1]$, and the square root operates on the elements of the initial error covariance matrix $\bm{P}_{0}$, which is defined as:
$\bm{P}_{0}=\textrm{diag}(\sigma_{\textrm{ r}}^{2}\bm{I}_{3x3},\sigma_{\textrm{v}}^{2}\bm{I}_{3x3},\sigma_{\textrm{R}}^{2}\bm{I}_{3x3},\sigma_{\textrm{SRP}}^{2}\bm{I}_{3x3})$ (46)
The performance analyzed hereafter is relative to case 1 outlined in Sec. 4.2. Figures 7 and 8 show the position and velocity error profiles and $3\sigma$ covariance bounds in the J2000 ecliptic reference frame on the studied trajectory leg. The sample error profile is displayed with blue solid lines, whereas the orange solid lines and the dashed ones define the $3\sigma$ covariance bounds of the samples and the filter, respectively.
Figure 7: Estimated errors for each position component with related $3\sigma$ bounds.
@@
  • Case 2: when the light-aberration effect on the planet position is corrected inside the IP (Sec. 3.2).
  • Case 3: when both effects are not corrected
  • Case 4: when only the light-aberration effect is compensated
  • Case 5: when only the light-time effect is compensated
@@
The root-mean-square error (RMSE) is chosen as the performance index to measure the estimation accuracy of the different models for the selected dataset. The latter comprises 50 initial state vectors and 814 images for simulation representing the scenarios observed during the interplanetary transfer. The filter and camera settings are unchanged with respect to case 1 for the sake of comparison.
Figure 9: Position and Velocity RMSE for the five cases.
@@
The estimation is biased when the corrections of the light effects are not taken into account in the filter (case 3). Between the two effects, the greater deviation from the nominal value is obtained when the light-time effect is not compensated (case 4). It can be noted that the RMSE associated with case 4 is greater than the one of case 3. This means that the deviation caused by the light-time effect is partially compensated by the light aberration. When the light-aberration correction is not applied on both the stars and planets centroids, the probe state estimation is affected by a small deviation (case 5). This occurs because the measurement model and the external observation are both affected by the light-aberration effect. Indeed, on one side, the attitude adopted to evaluate the measurement model is determined from stars in the image whose positions are aberrated. So, the measurement model is perturbed as well. On the other side, the external observation, i.e., the planet centroids, is not corrected for the light-aberration effect. Thus, the observation and the model are coherent.
Whereas, the performance of the filter when the light effects are both compensated in the measurement model (case 1) is analogous to the performance of the filter when the light aberration is corrected in the IP procedure (case 2). A slight discrepancy exists since the methodology followed to compensate for the light effects is different. In case 2, the light-aberration effect on the planet position is taken into account by correcting the observation, instead, in case 1, the light-aberration effect is compensated in the measurement model. Moreover, in case 1, the light-aberration effect is also taken into account in the derivation of $\bm{H}$ (Eq. 42). Nevertheless, the similarity of the performance allows us to validate the innovative measurement model proposed in Sec. 4.2 that provides an analytical first-order approximation of light effects corrections.
@@ 6 Conclusion
This paper develops an autonomous vision-based navigation algorithm for interplanetary transfer with an application for CubeSats missions. A non-dimensional extended Kalman filter adopting the planet position projection as external observation is chosen by considering the limited processing capabilities of a standard miniaturized processor. Moreover, the measurements exploited for the estimation correction are directly extracted from images generated with a deep-space rendering engine. This procedure allows for obtaining a more faithful value of the measurement error and its influence on the filter solution. At the end of the Earth–Mars trajectory, the filter estimates the spacecraft position and velocity with an accuracy of 2000 km and 0.5 m/s, respectively. Future analysis should examine the performance of the filter over low-thrust trajectories, which are desirable in CubeSats applications. Moreover, in this work, the probe attitude is determined from deep-space images with state-of-the-art star identification and attitude determination algorithms. The integration of the attitude filter in the proposed vision-based navigation algorithm and the adoption of a more robust star identification technique are objects of future investigation. Eventually, to validate the vision-based navigation applicability, it will be tested through hardware-in-the-loop simulations. Preliminary results have been already obtained through, first, the validation of the orbit determination algorithm with processor-in-the-loop simulations [9] and, second, with the validation of the image processing pipeline with the optical facility in the loop [45, 46].